Problem 2 Imagine a database from the Gallogly College of Engineering that contained the starting salaries for a sample of Spring 2018 graduating engineers. Recreate the following table and fill out at the intersections what happens to each of the sample statistics when every salary is altered by the specified amount. Due to an increase in every salary by Change in Mean Median Standard deviation $1000 5%
Answer #1

for increment of 1000:

new value X =Y+1000

mean E(X)=E(Y)+1000

medium =median(Y)+1000

std deviation SD(X)=SD(Y)

for increment in 5%

X=1.05Y

mean E(X)=1.05*E(Y)

median =1.05/median(Y)

std deviation =1.05*SD(Y)

therefore :

change in increase of $1000 increase of 5%
mean increase by 1000 increase by 5%
median increase by 1000 increase by 5%
std deviation does not change increase by 5%